# Nuclear level density of 128Te from $(\mathrm{p},\mathrm{p}^{\prime}\gamma)$ scattering and complementary photonuclear data

### v1

@@ 5 Conclusion and outlook
We have extracted the nuclear level density of 128Te from a $(\mathrm{p},\mathrm{p}^{\prime}\gamma)$ scattering experiment using the large-volume LaBr3:Ce and CeBr3 detectors from ELI-NP at the 9 MV tandem facilities at IFIN-HH. The decay data were normalised using photonuclear data, which provides nuclear level densities without intrinsic model dependencies from the constant temperature or Fermi gas models. The nuclear level density follows closely between the expectations from these two models, but we observe a clear divergence from the microscopic model based on the Skyrme force. Further work will include the in-depth study of the $\gamma$-ray strength functions from this data, in particular, to investigate how assumptions like the validity of the Brink-Axel hypothesis and the contribution from the spin distribution affect $\gamma$-ray strength-function results in both photonuclear data and charged particle data.
@@ Acknowledgements
PAS, AK, RB, MB, CCo, NMF, AGa, RL, CMi, AS, DAT, AT, GVT, and SU were supported by the ELI-RO program funded by the Institute of Atomic Physics, Măgurele, Romania, contract number ELI-RO/RDI/2024-002 (CIPHERS) and SA, DLB, SRB, RC, and TP ELI-RO/RDI/2024-007 (ELITE). The remaining authors from ELI-NP acknowledge the support of the Romanian Ministry of Research and Innovation under research contract PN 23 21 01 06. This project has received funding from the European Union’s Horizon Europe Research and Innovation programme under Grant Agreement No 101057511 (EURO-LABS).

### v2

@@ 5 Conclusion and outlook
We have extracted the nuclear level density of 128Te from a $(\mathrm{p},\mathrm{p}^{\prime}\gamma)$ scattering experiment using the large-volume LaBr3:Ce and CeBr3 detectors from ELI-NP at the 9 MV tandem facilities at IFIN-HH. The decay data were normalised using the photoabsorption cross-section, which provides nuclear level densities without intrinsic model dependencies from the constant temperature or Fermi gas models. The nuclear level density follows in between the expectations from these two models, but we observe a clear divergence from the microscopic model based on the Skyrme force. Further work will include the in-depth study of the $\gamma$-ray strength functions from this data, in particular, to investigate how assumptions like the validity of the Brink-Axel hypothesis and the contribution from the spin distribution affect $\gamma$-ray strength-function results in both photonuclear data and charged particle data. While the work presented here aimed at removing some of the model dependencies present in the analysis methods, there are still uncertainties present regarding, for example, the spin distribution that is difficult to resolve completely with current data and methods. Systematic investigations of pure $J^{\pi}=1^{-}$ NLDs with the future ELI-NP photon beams will further help pin down these uncertainties and the modelling of NLDs, partially addressing the remaining uncertainties or method dependent characteristics of NLD data.
@@ Acknowledgements
PAS, AK, RB, MB, CCo, NMF, AGa, RL, CMi, AS, DAT, AT, GVT, and SU were supported by the ELI-RO program funded by the Institute of Atomic Physics, Măgurele, Romania, contract number ELI-RO/RDI/2024-002 (CIPHERS) and SA, DLB, SRB, RC, and TP ELI-RO/RDI/2024-007 (ELITE). The remaining authors from ELI-NP acknowledge the support of the Romanian Ministry of Research and Innovation under research contract PN 23 21 01 06. This project has received funding from the European Union’s Horizon Europe Research and Innovation Programme under Grant Agreement No 101057511 (EURO-LABS). We would also like to acknowledge the International Atomic Energy Agency Coordinated Research Activity F41034 “Updating and Improving Nuclear Level Densities for Applications”. This work was presented at the International Symposium on Nuclear Science (ISNS-24), Sofia 2024.